Five kilograms of steam occupies a volume of 10m^3. Find the quality and the pressure if the temperature is measured at (a) 40o​C and (b) 86​0C.

ANS: (a) 0.1024, 7.383 kPa (b) 0.7312, 60.3 kPa
1
Expert's answer
2020-07-29T06:45:25-0400

Volume of steam(Vs)=10m^3,

Mass of steam(Ms)=5 Kg

Therefore specific volume of steam(v)=10/5=2 m^3/Kg

(a) From Steam table: At Temperature(T)=40C, Saturation pressure(P)=7.3849 kPa.

vf=0.00100789 m^3/kg, vg= 19.515 m^3/kg

v=vf+xvfg

2=0.00100789+x(19.51399)

Quality of steam(x)=0.1024

(b) From Steam table: At Temperature(T)=86C, Saturation pressure(P)=60.173 kPa.

vf=0.00103312 m^3/kg, vg= 2.7244 m^3/kg, vfg= 2.723366 m^3/kg

v=vf+xvfg

2=0.00103312 +x( 2.723366 )

Quality of steam(x)=0.7340